I need help on this word problem. I am not sure what method to use and how to solve it:
Mike made a trip in 40 hours. Joe made the same trip in 30 hours because Joe was traveling 6km per hour faster. How many kilometers was the trip?

Answer by scott8148(5891)

d = r * t, so r = d / t

"Mike made a trip in 40 hours" __ r = d / 40

"Joe made the same trip in 30 hours because Joe was traveling 6km per hour faster" __ r + 6 = d / 30

substituting __ (d/40)+6 = d/30

multiplying by 120 (LCD) __ 3d + 720 = 4d

subtracting 3d __ 720 = d